A rocket of mass 20 kg has 180 kg of fuel. The exhaust velocity of fuel is 1.6km/sec. Calculate the ultimate velocity of the rocket gained ,when the rate if consumption if the fuel is 2kg/sec (neglect gravity)

Grade:10

3 Answers

Vikas TU
4 years ago
Dear student 
Velocity of rocket is given as

v=u ln(M/M-rt)

where M is the total mass of rocket + fuel at t=0 sec = 20 +180 =200 kg

and r rate of fuel consumption =2kg/sec

u is the velocity of gas ejection w.r.t rocket = 1.6 km/sec

ultimate velocity is achieved when all the fuel will burn

ie. M-rt =20 kg

so v= 1.6 ln(180/20)
=3.515 km/sec
Arun
4 years ago
 

Velocity of rocket is given as 

v=u ln(M/M-rt)

wher M is the total mass of rocket + fuel at t=0 sec = 20 +180 =200 kg

and r rate of fuel consumption =2kg/sec

u is the velocity of gas ejection w.r.t rocket = 1.6 km/sec

 

ultimate velocity is achieved when all the fuel will burn 

ie. M-rt =20 kg

so v= 1.6 ln(180/20)

        =3.515 km/sec

aswanth nayak
4 years ago
Dear Student,
 
Velocity of rocket is given as, v=u ln(M/M-rt)

where M is the total mass of rocket + fuel at t=0 sec = 20 +180 =200 kg

and r rate of fuel consumption =2kg/sec

u is the velocity of gas ejection w.r.t rocket = 1.6 km/sec

ultimate velocity is achieved when all the fuel will burn

ie. M-rt =20 kg

so v= 1.6 ln(180/20)
==>3.515 km/sec
 
Hope it helps you
Regards